OpenStudy (anonymous):

write 3(cos225 +isin225) in the form a + ib.

OpenStudy (zehanz):

225 degrees rings a bell, no?

OpenStudy (anonymous):

it is complex numbers

OpenStudy (zehanz):

Not yet ;) I meant that sin225 and cos225 are supposed to be well-known numbers: they have to do with the 45-45-90 degree triangle, with sides a, a and a√2. Does that ring a bell? See image. Both sin225 and cos225 are: \[-\frac{ 1 }{ 2 }\sqrt{2}\]

OpenStudy (anonymous):

you mean right angled triangle?

OpenStudy (zehanz):

yeah

OpenStudy (anonymous):

so you mean the angles should be in the first quadrant?

OpenStudy (zehanz):

No, 225 is in the third quadrant. I only meant that in the 3rd quadrant you get the same values (just negative) as in the triangle with 45-45-90 degrees.

OpenStudy (zehanz):

Therefore you now have \[3(-\frac{ 1 }{ 2 }\sqrt2+i \cdot -\frac{ 1 }{ 2 }\sqrt 2)\]And if you work out the brackets you're done.

OpenStudy (anonymous):

Yeah I get you but that does not make any change.

OpenStudy (zehanz):

Is your problem solved then?